1. Pyridine (C5H5N) is a particularly foul-smelling substance used in manufacturing pesticides and plastic resins. Calculate the pH of a 0.145 M solution of pyridine (Kb = 1.4×10-9 ).
1
Expert's answer
2020-04-13T02:24:58-0400

C5H5N +H2O <=>HC5H5N+ + OH-

Kb=[HC5H5N+][OH-] / [C5H5N]=1.4*10-9

Let [HC5H5N+]=[OH-]= x, then [C5H5N]=0.145-x

1.4*10-9=x*x/0.145-x;

1.4*10-9*(0.145-x)=x2

x2 +1.4*10^-9*x - 2.03*10^-10=0;

x has no solution